# Universal neighborhood topology and geometry of exceptional points in physical systems

### v1

@@ Abstract
Exceptional points (EPs) occurring in non-Hermitian systems at certain physical parameters are intensively studied in many areas of physics, including diffraction optics, lasers, atomic and polaritonic condensates, often in the context of sensing. Recent discoveries of EPs in nonlinear systems open the door for an even larger parameter space, raising the question of whether the geometric structure of EPs is universal and independent of the physical model. We show that this is the case for nonlinear perturbations of an isolated 2nd-order linear EP which becomes the organizing point of a universal elementary catastrophe (elliptic umbilic). This clarifies not only the neighborhood’s topology but also its geometric shape (cone with quasi-deltoid cross section). Thus, the position and characteristics of EPs can be predicted in nonlinear non-Hermitian parameter space; e.g., at a 2nd-order linear EP four nonlinear eigenvectors coalesce. These fundamental insights on universal topological structures and phase boundaries accompanying EPs in nonlinear physical systems will pave the way for the purposeful design of such systems with novel functionalities and control possibilities.
Over the last decade or so, the physics of exceptional points (EPs), which are commonly defined as non-Hermitian degeneracies with coalescing eigenvalues and eigenvectors (mathematical definitions and generalizations are in kato.1966 ; ashida-etal.2021 ), has been increasingly explored across a broad range of scientific areas, including anharmonic oscillators bender-wu.1969 , diffraction optics berry-odell.1998 , quantum chaos leyvraz-heiss.2005 ; heiss.2012 , conventional lasers brandstetter-etal.2014 ; lietzer-etal.2012 , waveguides and photonic systems kullig-etal.2023 ; longhi.2018 ; quirozjuarez.2019 ; dong-etal.2020 , atomic gases, condensates and high-Q cavity atomic lasers xu-etal.2017 ; choi-etal.10 ; wang-etal.2024 , polariton lasers, microlasers and polariton condensates dembowski-etal.01 ; miao-etal.2016 ; gao-etal.15 ; gao2018chiral ; ozturk-etal.2021 ; khurgin.2020 ; Li2022 ; PhysRevLett.122.185301 ; PhysRevB.104.235408 ; Li2022 , non-Hermitian Bose–Hubbard models graefe-etal.2008 , critical fluctuations and fluctuation spectra hanai-littlewood.20 ; binder-kwong.2021 , and sensing applications wiersig.2014 ; hodaei-etal.2017 ; langbein.2018 ; wiersig.2020 ; sakhdari-etal.2019 , non-Hermitian parity-time (PT) symmetric quantum systems bender-hook.2024 ; hassan-etal.2015 ; miri-alu.2016 ; ge-elganainy.2016 ; kominis-etal.2017 ; teimourpour-etal.2017 ; xia-etal.2021 ; lee-etal.2024 ; klauck-etal.2025 , and quantum state control abbasi-etal.2022 . EPs have also been studied in the context of topological insulators and other topological systems (e.g. luitz-etal.2019 ; kawabata-etal.2019 ; tang-etal.2020 ; bergholtz-etal.2021 ; hu-etal.2022 ; yokomizo-etal.2020 ; ding-etal.2022 ). While in most of these physical realizations EPs emerge as eigenvalues $\lambda$ of a linear non-Hermitian matrix $M$, i.e. from the eigenvalue problem $M\textbf{x}=\lambda\textbf{x}$, more recent studies of EPs include nonlinear physical systems, such as optical or electric resonators, as well as polariton condensates. In these systems, the complex interplay of nonlinear and non-Hermitian physics shifts PhysRevA.103.043510 ; el2023tracking ; Ramezanpour:24 ; wingenbach-etal.2024 and rotates wingenbach-etal.2024 the EP in parameter space by blueshifts and saturable gain. Moreover, mode switching can be realized in bistable regions near the EP wang2019dynamics . In Refs. wingenbach-etal.2024 ; Li2024 , it was shown that nonlinearity can alter the eigenvalue splitting in the vicinity of an EP and highly sensitive EP-based sensors have been realized in nonlinear non-Hermitian systems PhysRevApplied.18.054059 ; Bai-etal.2023 .
The eigenvalue problem may be nonlinear in the sense that the matrix $M$ may depend on the eigenvector x. For example, the energies in an optical microcavity, which are described by the diagonal elements of $M$, can shift with the intensity of the light field described by x, hence $M(\textbf{x})\textbf{x}=\lambda\textbf{x}$. In the following, we refer to EPs that are solutions of a linear eigenvalue problem as ‘linear EPs,’ and those solving a nonlinear eigenvalue problem are called ‘nonlinear EPs.’
@@
The wide range of physical systems in which EPs can occur raises the question whether there exist any guiding principles that help clarify the occurrence of EPs as the physical parameters entering the matrix $M$ in each of the systems are varied, for example, the mode coupling in classical diffraction gratings berry-odell.1998 or frequency and coupling constants of ring resonators in sensing applications wiersig.2020 ).
@@
In this paper, we demonstrate that general insight can be obtained on the distribution of EPs in parameter space, especially for nonlinear eigenvalue problems. We show that a powerful way to do this is to consider the behavior of eigenvectors around EPs as a bifurcation problem and make use of established mathematical results in catastrophe theory. We illustrate this procedure with an analysis of (nonlinear) parameter-dependent $2\times 2$ matrices and show that, for a broad class of physically relevant nonlinear eigenvalue problems, there are general constraints to the topology and geometric shape of the set of nonlinear EPs. These constraints are universal and independent of the physical system that the matrix $M$ represents.
An EP occurs at specific parameter values. Let us assume the linear EP to be isolated, occurring at certain values of the parameters entering the linear matrix $M$. We can vary the parameters away from the EP, in other words look at the neighborhood of the EP, and, by assumption, we will not find any other linear EPs in a small neighborhood. But if we allow the matrix to be nonlinear, the question about EPs in the neighborhood becomes more complicated. To answer the question, we would have to solve the nonlinear eigenvalue problem in order to find the location of the EP – or possibly even a multitude of EPs. We could then see if, for a given nonlinearity, we have more EPs or if the EP goes away once the nonlinearity is switched on. But in order to get a complete picture of the EPs existing in the nonlinear case, we would have to fix the nonlinearity and then solve the nonlinear eigenvalue problem for all remaining parameters entering the matrix. It is a priory not even clear how many remaining parameters we have to take into consideration. For example, in the case of the coupled ring resonators, we could limit ourselves to cases without loss and with each resonator having the same frequency (3 parameters, cf. Eq. (3) in wiersig.2020 ), but we could also generalize the model if we assume each resonator to have a different frequency and different loss (5 parameters). And we would have to do this ambitious analysis for each fixed nonlinearity. Finally, we would have to carry out this program for each physical realization of EPs (diffraction optics, polariton physics, sensing applications, etc.) separately. This means that, realistically speaking, it is difficult to get a full understanding of the structure of the set of EPs.
@@
To solve the highly sophisticated problem outlined above, we utilize the general theorems of a branch of bifurcation theory called catastrophe theory zeeman.1976b ; thom.1989 ; arnold.1992 ; saunders.80 ; gilmore.81 ; poston-stewart.1978 . Roughly speaking, the framework of catastrophe theory describes the shapes of singularity sets – or phase boundaries in nonequilibrium phase diagrams, to use a physicist’s language – related to singularities of potential functions $V_{L}$ (Lyapunov potentials). The word catastrophe is used for the sudden qualitative change or bifurcation at a ‘phase boundary.’ Moreover, it lists all possible shapes (‘elementary catastrophes’) for cases in which the number of variables (often called state variables) and the number of control parameters (i.e. external parameters) are relatively small, see Supplemental Material (SM), Table 1).
@@
Examples of systems that have been analyzed by catastrophe theory include dynamical systems obeying equations of motion of the form $=-{\rm grad}\{V_{L}(\textbf{x})\}$ x ) }, where the phase boundaries in the control parameter space follow from the assumption of degenerate stationary states (degenerate critical points), corresponding to ${\rm grad}\{V_{L}(\textbf{x})\}=0$ and ${\rm det}\{h\}=0$, where $h$ denotes the Hessian matrix of $V_{L}$. Similarly, in optics, the solution of the Helmholtz equation can be evaluated with the method of stationary phase and stationary points, as well as caustics occuring when two or more stationary points coalesce (p. 336 of adam.2002 ). This identifies rainbows and other optical phenomena as diffraction catastrophe, specifically fold and cusp bifurcations; see berry-upstill.1980 ; berry.90 ; adam.2002 and Ch. 12 of poston-stewart.1978 for details. Higher bifurcations (catastrophes), such as the so-called elliptic umbilic, have also been studied in the context of optical diffraction berry-etal.79 ; berry.90 . Many other areas of application of catastrophe theory can be found in zeeman.1976a ; zeeman.1976b ; saunders.80 ; gilmore.81 ; rosen.1988 ; poston-stewart.1978 ; heiss.04 ; longhi.2018 ; heiss.2012 ; barkleyrosser.2007 . This further extends to studies in biology, social sciences, and economics, with various degrees of acceptance because low-dimensional dynamical-systems models are not always appropriate for the simulation of complex systems in those sciences rosen.1988 ; barkleyrosser.2007 . In contrast, in many areas of physics low-dimensional models and eigenvalue problems (e.g. two–level models) are well established.
@@
Figure 1: (a) Sketch of a non-Hermitian and nonlinear two-mode system, with mode coupling $\beta$, mode energy splitting $2\delta=\delta_{x}-\delta_{y}$, and loss/gain difference $2\gamma$. (b) Sketch of hypothetical location of exceptional points as a function of physical parameters. Without a nonlinearity, $\alpha=0$, isolated linear EPs are at $(\beta,\delta)=(\pm\gamma,0)$; in the nonlinear regime, $\alpha\neq 0$, EPs are implied on the colored surfaces. To illustrate the problem our study is solving, the surfaces we show here are one of infinitely many hypothetically possible (but generally wrong) EP neighborhoods. The correct geometric shape is shown in Fig. 2. Panels (c) and (d) show the eigenvalues in the linear case ($\alpha=0$) with two isolated EPs at $(\beta,\delta)=(\pm\gamma,0)$.
@@
Figure 2: Surface of EP locations as a function of physical parameters (singularity set) explained in Fig. 1. In the weakly nonlinear regime, $\alpha\ll\gamma$, the shape approaches that of an elliptic umbilic bifurcation (‘catastrophe’). The shape of each singularity set consists of two infinite three–cusped conical surfaces, the apices of which meet at the two linear EPs $(\alpha,\beta,\delta)=(0,\pm\gamma,0)$. At large $\alpha$, the three–cusped shape remains but is deformed guaranteeing that the two cone-like surfaces originating at $\beta=\pm 1$ do not overlap; see Fig. 3.
@@
Figure 3: Cross section of EP surface for fixed nonlinearity $\alpha$; axis scaled with scale factor $s=2(\sqrt{1+\alpha^{2}/\gamma^{2}}-1)$. Here, $\beta^{\prime}=\beta-\gamma$. Small $\alpha\ll\gamma$ (vicinity of linear EP, Eq. 5): deltoid or three-cusped hypocycloid broecker-lander.1975 ; poston-stewart.1978 ; saunders.80 of the elliptic umbilic surface with exact scaling $s=\alpha^{2}/\gamma^{2}$. Large $\alpha$ (farther away from linear EP, Eq. 4): exact scaling is lost, but three–cusped shape remains. The half-infinite line where PT symmetry is unbroken is indicated by the blue line.
In our study, the elements of the eigenvectors of the non-Hermitian eigenvalue problem $[M(\textbf{x})-\lambda]\textbf{x}=0$ are the state variables. Our strategy is then as follows: (more details are given in the SM): The eigenvalue equation constitutes two equations for the unkowns that we call state variables (and which determine the eigenvector elements; see below). We write those equations as gradients of an appropriately chosen Lyapunov potential, i.e. in the form ${\rm grad}\{V_{L}(\textbf{x})\}=0$. The condition for an EP (coalescing eigenvectors) is then the same as ${\rm det}\{h\}=0$. Finally, we perform a Taylor expansion of $V_{L}$ around the EP, allowing us to identify the Lyapunov potential with one of the elementary catastrophes, listed for example on p. 137 of saunders.80 and p. 11 of gilmore.81 , and reproduced in the SM. This then fixes the topological and geometric structure of the EP neighborhood. Below, we illustrate our approach for $2\times 2$ matrices with a second–order linear EP. In that case, we find that linear EPs are the organizing point of an elliptic umbilic singularity.
We write the nonlinear eigenvalue problem (now using the symbol $H$ as commonly used in the context of non-Hermitian Hamiltonians) as

### v3

@@ Abstract
Exceptional points (EPs) are non-Hermitian degeneracies where eigenvalues and eigenvectors coalesce, giving rise to unusual physical effects across scientific disciplines. The concept of EPs has recently been extended to nonlinear physical systems. We theoretically demonstrate a universal topology in the nonlinear parameter space for a large class of physical systems that support $2^{\mathrm{nd}}$ order EPs in the linear regime. Knowledge of this topology (called elliptic umbilic singularity in bifurcation theory) deepens our understanding of $2^{\mathrm{nd}}$ order linear EPs, which here emerge as coalescence of 4 nonlinear eigenvectors. This helps guide future experimental discovery of nonlinear EPs and their classification, establish rigorous bounds of sensitivity enhancement of EPs in nonlinear systems, and helps envision and optimize technological applications of nonlinear EPs. Our theoretical approach is general and can be extended to nonlinear perturbations of $3^{\mathrm{rd}}$ and higher-order EPs.
Exceptional points (EPs), which are spectral degeneracies of non-Hermitian systems at which both eigenvalues and eigenvectors coalesce (mathematical definitions and generalizations are in [1, 2, 3]), have been increasingly explored across a broad range of scientific areas, including anharmonic oscillators [4], diffraction optics [5], quantum chaos [6, 7], conventional lasers [8, 9], waveguides and photonic systems [10, 11, 12, 13], atomic gases, condensates and high-Q cavity atomic lasers [14, 15, 16], polariton lasers, microlasers and polariton condensates [17, 18, 19, 20, 21, 22, 23, 24, 25, 23], non-Hermitian Bose–Hubbard models [26], critical fluctuations and fluctuation spectra [27, 28], and sensing applications [29, 30, 31, 32, 33, 34], non-Hermitian parity-time (PT) symmetric quantum systems [35, 36, 37, 38, 39, 40, 41, 42, 43, 44], and quantum state control and squeezing [45, 46]. EPs have also been studied in the context of topological insulators and other topological systems (e.g. [47, 48, 49, 50, 51, 52, 53, 54, 55, 56]).
While in most of these physical realizations EPs emerge as eigenvalues $\lambda$ of a linear non-Hermitian matrix $M$, i.e. from the eigenvalue problem $M\textbf{x}=\lambda\textbf{x}$, more recent studies of EPs include nonlinear physical systems, such as optical or electric resonators, [57], models of population dynamics [58], as well as polariton condensates. In these systems, the complex interplay of nonlinear and non-Hermitian physics leads to counterintuitive phenomena such as shifts [59, 60, 61, 62] and rotations [62] of the EP in parameter space through blueshifts and saturable gain. Moreover, mode switching can be realized in bistable regions near the EP [63]. In Refs. [62, 64], it was shown that nonlinearity can alter the eigenvalue splitting in the vicinity of an EP and highly sensitive EP-based sensors have been realized in nonlinear non-Hermitian systems [65, 66, 67]. The underlying eigenvalue problem here may be nonlinear in the sense that the matrix $M$ may depend on the eigenvalue $\lambda$ (e.g. [68]), the eigenvector x (e.g. [69]), or both. In some instances, the nonlinear dependence on the eigenvector can be interchanged by that on the eigenvalue, e.g. [34]. Other kinds of nonlinear eigenvalue problems are discussed in [70]. In the following, we concentrate on the case where the nonlinear eigenvalue problem is formulated as a matrix equation where the matrix dependends on the eigenvector. For example, the energies in an optical microcavity, which are described by the diagonal elements of $M$, can shift with the intensity of the light field described by x, hence $M(\textbf{x})\textbf{x}=\lambda\textbf{x}$. In the following, we refer to EPs that are solutions of a linear eigenvalue problem as ‘linear EPs,’ and those solving a nonlinear eigenvalue problem are called ‘nonlinear EPs.’
@@
The wide range of physical systems in which EPs can occur raises the question whether there exist any guiding principles that help clarify the occurrence of EPs as the physical parameters (in particular the nonlinearity) entering the matrix $M$
are varied.
@@
In this paper, we demonstrate that important insight can be obtained on the distribution of EPs in parameter space, especially for nonlinear eigenvalue problems. We show that a powerful way to do this is to consider the behavior of eigenvectors around EPs as a bifurcation problem and make use of established mathematical results in catastrophe theory.
We illustrate this rigorous procedure with an analysis of nonlinear parameter-dependent $2\times 2$ matrices, schematically shown in Fig. 1, and show that a broad class of physically different nonlinear systems share the same EP neighborhood geometry that fixes one (and only one) topology, namely a cone-like structure (with deltoid cross section) that consists of three fold-bifurcation surfaces attached to each other at cusp-bifurcation ribs. As a corollary, we find the possible numbers of solutions (eigenvalues) in each region of parameter space. All these insights are based on one single equation (namely that for the so-called Lyapunov potential). Our findings can be used in future studies to establish rigorous bounds on the sensitivity enhancement of EPs in nonlinear systems. It also provides guardrails against possibly wrong conclusions, illustrated in Fig. 1(b), that might be obtained from limited experimental observations or approximate theories.
@@
To solve the highly sophisticated problem outlined above, we utilize the general theorems of a branch of bifurcation theory called catastrophe theory [71, 72, 73, 74, 75, 76]. Roughly speaking, the framework of catastrophe theory describes the shapes of singularity sets – or phase boundaries in nonequilibrium phase diagrams, to use a physicist’s language – related to singularities of potential functions $V_{L}$ (Lyapunov potentials). The word catastrophe is used for the sudden qualitative change or bifurcation at a ‘phase boundary.’ Moreover, it lists all possible shapes (‘elementary catastrophes’) for cases in which the number of variables (often called state variables) and the number of control parameters (i.e. external parameters) are relatively small, see Supplementary Table 1 of the Supplementary Information.
@@
Examples of systems that have been analyzed by catastrophe theory include dynamical systems obeying equations of motion of the form $\dot{\textbf{x}}=-{\rm grad}\{V_{L}(\textbf{x})\}$, where the phase boundaries in the control parameter space follow from the assumption of degenerate stationary states (degenerate critical points), corresponding to ${\rm grad}\{V_{L}(\textbf{x})\}=0$ and ${\rm det}\{h\}=0$, where $h$ denotes the Hessian matrix of $V_{L}$. Similarly, in optics, the solution of the Helmholtz equation, which describes the stationary solutions of a second-order time-differential equation, in contrast to the example of a first-order time-differential equation mentioned above, can be evaluated with the method of stationary phase and stationary points, as well as caustics occurring when two or more stationary points coalesce (p. 336 of [77]). This identifies rainbows and other optical phenomena as diffraction catastrophe, specifically fold and cusp bifurcations; see [78, 79, 77] and Ch. 12 of [76] for details. Higher bifurcations (catastrophes), such as the so-called elliptic umbilic, have also been studied in the context of optical diffraction [80, 79]. Many other areas of application of catastrophe theory can be found in [81, 71, 74, 75, 82, 76, 83, 11, 7, 84]. This further extends to studies in biology, social sciences, and economics, with various degrees of acceptance because low-dimensional dynamical-systems models are not always appropriate for the simulation of complex systems in those sciences [82, 84]. In contrast, in many areas of physics low-dimensional models and eigenvalue problems (e.g. two–level models) are well established.
@@
Figure 1: Non-Hermitian two-state system. (a) Sketch of a non-Hermitian and nonlinear two-state (or two-mode) system, with linear coupling $\beta$, energy splitting $2\delta=\delta_{x}-\delta_{y}$, and loss/gain difference $2\gamma$. (b) Sketch of hypothetical location of exceptional points as a function of physical parameters. Without a nonlinearity, $\alpha=0$, isolated linear EPs are at $(\beta,\delta)=(\pm\gamma,0)$; in the nonlinear regime, $\alpha\neq 0$, EPs are implied on the colored surfaces. To illustrate the problem our study is solving, the surfaces we show here are one of infinitely many hypothetically possible (but generally wrong) EP neighborhoods. The correct geometric shape is shown in Fig. 2. Panels (c) and (d) show the eigenvalues in the linear case ($\alpha=0$) with two isolated EPs at $(\beta,\delta)=(\pm\gamma,0)$.
@@
Figure 2: Elliptic umbilic singularity set. Surface of EP locations as a function of physical parameters (singularity set) explained in Fig. 1. In the weakly nonlinear regime, $\alpha\ll\gamma$, the shape approaches that of an elliptic umbilic bifurcation (‘catastrophe’). The shape of each singularity set consists of two infinite three–cusped conical surfaces, the apices of which meet at the two linear EPs $(\alpha,\beta,\delta)=(0,\pm\gamma,0)$. At large $\alpha$, the three–cusped shape remains but is deformed guaranteeing that the two cone-like surfaces originating at $\beta=\pm 1$ do not overlap; a see Fig. 3.
@@
Figure 3: Cross section of singularity set. Cross section of EP surface for fixed nonlinearity $\alpha$; axis scaled with scale factor $s=2(\sqrt{1+\alpha^{2}/\gamma^{2}}-1)$. Here, $\beta^{\prime}=\beta-\gamma$. Small $\alpha\ll\gamma$ (vicinity of linear EP, Eq. 6): deltoid or three-cusped hypocycloid [85, 76, 74] of the elliptic umbilic surface with exact scaling $s=\alpha^{2}/\gamma^{2}$. Large $\alpha$ (farther away from linear EP, Eq. 5): exact scaling is lost, but three–cusped shape remains. The half-infinite line where PT symmetry is unbroken is indicated by the blue line.
In this study, we aim to understand the topological structure of the EP set around the linear EP in parameter space when nonlinearity in the matrix is turned on. Our strategy of utilizing catastrophe theory is as follows (more details are given in Methods and Sec. II of the Supplementary Information). The layout of the catastrophe theory formulation consists of a state variable space and a control parameter space. The eigenvector elements of the nonlinear matrix are considered as values of state variables determined by the eigenvector equations, and the coordinates of the control parameter space are the parameters in the matrix. A Lyapunov potential function is constructed under the condition that its critical points (in state variable space) are equal to the eigenvector elements. The coalescing eigenvectors are accordingly given by the degenerate critical points of the Lyapunov potential. The set of degenerate critical points forms a hyper-surface in the combined state variable space and parameter space. The projection of this hyper-surface onto the parameter space yields the set of EPs in parameter space.
To obtain a topological characterization of the nonlinear EP set in parameter space around the linear EP, we Taylor expand the Lyapunov potential in both the state variables and the parameters about this point. This allows us to match our Lyapunov potential locally with one of the known catastrophes. See, for example, p. 137 of [74] and p. 11 of [75] for a list of the ‘elementary catastrophes’, which is also reproduced as Supplementary Table I in the Supplementary Information. We find that, for the broad class of nonlinear $2\times 2$ matrices considered here, the matched catastrophe is the elliptic umbilic. Consequently, the EP sets (in parameter space) of these matrices share a common topological structure which is that of the set of degenerate critical points of this catastrophe.
1 Results
In the following, we present our approach for a large class of $2\times 2$ matrices and analyze the EP structure in parameter space. We find that, for all matrices in this class, the linear EPs are the organizing points of an elliptic umbilic singularity. We stress that, in addition to the matrices discussed below, there may be other $2\times 2$ matrices with elliptic umbilic nonlinear EP structures that we have not yet considered.
In the next subsection, Sec. 1.1, we illustrate our findings using a specific example, namely a symmetric matrix with Kerr-like nonlinearity, and in Sec. 1.2-1.4 we discuss extensions and generalizations of that example. All examples discussed below belong to the class of matrices that can be written in the following form (now using the symbol $H$ as commonly used in the context of non-Hermitian Hamiltonians)
$H=\begin{pmatrix}\delta+i\gamma+\alpha F(w)&b\\ c&-\delta-i\gamma-\alpha F(w)\end{pmatrix}\,$ (1)
with $\delta,\gamma,\in\mathbb{R}$, $b,c,\alpha\in\mathbb{C}$, and the nonlinearity $F(w)$ being a real-valued analytic function of the eigenvector parameter $w$ in the vicinity of the linear EP. Specifically, the real-valued variable $w$ is related to the (right) eigenvector of $H$, which, normalized to $1$, is written as $x\,\,y]^{T}=(1/\textit{A})[1\,\,\tilde{w}]^{T}$, $\tilde{w}=we^{i\theta}$, $\textit{A}=\sqrt{1+w^{2}}$. The matrix in Eq. (1) is a useful starting point for our analysis. The fact that $H$ is traceless, in other words that the average of the diagonal elements is zero, is not a restriction. Below, in Sec. 1.2, we will discuss in more detail the procedure of how to remove the average of the diagonal elements without changing the eigenvectors.
1.1 Example: quadratic nonlinearity
In order to give a concrete example, in this section we illustrate our approach for symmetric $2\times 2$ matrices with a second–order linear EP and a nonlinearity that is a quadratic function of the eigenvector variable $w$ defined above (Kerr-like nonlinearity). Furthermore, we assume that at the EP, where $w=1$, the lowest-order Taylor expansion of the nonlinearity function yields $F(1)=0$ and $F^{\prime}(1)\neq 0$. We write the nonlinear eigenvalue problem as